[figure: 9.7 cm]

THEORETICAL SEARCH FOR MAGNETIC HALF-HEUSLER SEMICONDUCTOR

J. Toboła^{a}, S. Kaprzyk^{a}, P. Pecheur^{b} aFaculty of Physics and Nuclear Techniques University of Mining and Metallurgy, 30-059 Kraków, Poland b Laboratoire de Physique des Materiaux, Ecole des Mines Parc du Saurupt, 54042 Nancy Cedex, France

We present electronic structure of half-Heusler CoTi_{1−x}T_{x}Sb and FeV_{1−x}T_{x}Sb with x = 0.001, 0.005, 0.01 and 0.05 derived from the charge and spin-selfconsistent KKR-CPA computations. Experimentally, both parent compounds were found to show semiconducting properties [1]. This was supported by the band structure calculations [2], which gives an energy gap as large as 0.95 eV (CoTiSb) and 0.46 eV (FeVSb). Among transition metal impurities doped in the aforementioned semiconductors, Cr and Mn atoms exhibit a strong polarisation of density of states, if substituting selectively on Tiand V-site in CoTiSb and FeVSb, respectively. The calculated magnetic moment on Mn is found to be on average as large as 2.5 μ_{B} and 3.2 μ_{B} in FeVSb and CoTiSb, respectively. Inspecting total DOS near the gap in these magnetic systems we find a behavior very similar of half-metallic magnetics than of magnetic semiconductors. Moreover, if impurity concentration decreases the energy gap tends to open for both spin channels.

[1] M. A. Kouacou et al., J. Phys. Condens. Matter 7 (1995) 7373; D. P. Young et al., J. Appl. Phys. 87 (2000) 317; Y. Xia et al., J. Phys. Condens. Matter 13 (2001) 77.

[2] J. Toboła et al. J. Alloys Compd. 296 (2000) 243.

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : oral

Corresponding author : J. Toboła

Address for correspondence :

1. Faculty of Physics and Nuclear Techniques, AGH Al. Mickiewicza 30, 30-059 Kraków, Poland

2. Laboratoire de Physique des Materiaux, Ecole des Mines Parc du Saurupt, 54042 Nancy Cedex, France

Email address : tobola@ftj.agh.edu.pl